# Detecting and Rectifying Noisy Labels: A Similarity-based Approach

### v1

@@ Abstract
Label noise in datasets could damage the performance of neural net training. As the size of modern deep networks grows, there is a growing demand for automated tools for detecting such errors. In this paper, we propose post-hoc, model-agnostic error detection and rectification methods utilizing the penultimate feature from a neural network. Our idea is based on the observation that the similarity between the penultimate feature of a mislabeled data point and its true class data points is higher than that for data points from other classes, making the probability of label occurrence within a tight, similar cluster informative for detecting and rectifying errors. Extensive experiments show our method not only demonstrates high performance across various noises but also automatically rectifies these errors to improve the quality of datasets [^1] Our implementation is available at https://github.com/huutiendang/noise-detection-and-rectification.
1 Introduction
@@
While the majority of knowledge in AI systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f. the LeCun’s cake). For Large Language Models (LLMs) such as GPTs Brown et al. (2020), LLaMA Touvron et al. (2023a, b), and Gemini Team et al. (2023), supervised learning accounts for only a small fraction of the total computation budget but has a significant impact on the models’ performance. Recent research (e.g. Zhou et al., 2023; Gunasekar et al., 2023) finds that high quality training data significantly improves performance while reducing the training cost by orders of magnitude. The need for automated tools for improving the quality of supervised learning data is rising as datasets and models are getting larger at an unprecedented speed.
Real-world datasets contain a notable amount of errors Beyer et al. (2020); Northcutt et al. (2021b). Previous works Dau et al. (2022); Nguyen-Duc et al. (2023) showed that removing errors from the training set improves the performance of AI models trained on that dataset. Automatic error rectification, however, is an underexplored topic. In this paper, we present a similarity based approach for error detection and rectification in large scale datasets. We provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Sec. 3.2). Inspired by this observation, we develop simple yet effective similarity-based methods for detecting and rectifying label errors (Sec. 3.3). Extensive experiments demonstrate the superiority of our methods across various settings (Sec. 4.2). Furthermore, our methods are posthoc and model-agnostic i.e., they can be applied to any deep neural network (DNN) architectures without the need for retraining.
2 Background and Related Work
@@
Confident-based approach.
@@
Confident-based methods are based on the notion of confident learning Northcutt et al. (2021a) that deriving label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al. (2022). Low confidence serves as a heuristic indicating the likelihood of label noise. Given a data point $\mathbf{z}$ with output label $\mathbf{y}=(y_{1},...,y_{k},...y_{N})$, the model’s predicted probabilities is $\mathbf{p}=(p_{1},...,p_{k},...,p_{N})$ over $N$ classes. Northcutt et al. (2021a) proposed three label quality scoring methods:
@@
(1) Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$: $\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}}$, for $k\in\{1,2,...,N\}$.
@@
(2) Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class: $\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}}$, for $j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$
(3) Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy: $\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{\mathcal{H}_{N}(\mathbf{p})}$, where $\mathcal{H}_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$.
Gradient-based approach.
@@
Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN. Dau et al. (2022) proposed a way to adapt IF and its variants i.e., Gradient Dot Product (GD; Charpiat et al. (2019a)), Gradient Cosine (GC; Charpiat et al. (2019a)), and Tracing Gradient Decent (TracIn; Pruthi et al. (2020)), for identifying errors in large-scale source code datasets. The idea is that the gradients of error data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points. The algorithm computes the influence score of each data point in the noisy dataset with data points in a reference set. A more negative influence score means it is more likely to be an error. Nguyen-Duc et al. (2023) use class information to improve the performance and stability of these gradient methods.
(1) $\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=-\frac{1}{n}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\mathcal{H}_{\hat{\theta}}^{-1}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})$,
@@
(2) $\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\langle\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})\rangle$,
(3) $\textnormal{GC}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\textnormal{cos}(\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)}))$,
(4) $\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\sum_{t=1}^{T}\eta_{t}\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$, where $\mathcal{H}_{\hat{\theta}}$ is the hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ is the learning rate at epoch $t$.
Other approaches.
The rule-based approach Chu et al. (2013) and statistics-based approach Huang and He (2018) are commonly used for structured data such as tabular data. These methods are not suitable for deep learning, as they assume convexity in the model, and the rules in many large-scale datasets are not easy to find and describe.
3 Method
3.1 Observation
We design experiments to randomly corrupt and inject noise into datasets. We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points. As an illustrative example in Fig. 2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points.
@@ 3.1 Observation
(a) Distribution of (a) Cosine similarity and (b) Dot product over IMDB Maas et al. (2011) with $10\%$ noise. Blue bars represent the similarity between mislabeled data points and their true class data points, red bars represent the similarity between mislabeled data points and other class data points. Features are obtained from a trained BERT model.
We find that this phenomenon persists across varying percentages of noise. To complement our observation, we provide a theoretical analysis in Section 3.2.
3.2 Theoretical Analysis
@@
We consider a deep network receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension. Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias. Given an output function $\phi_{W}$ parameterized by $W\in\mathbb{R}^{N\times d}$. For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we have two output vectors $\phi_{W}(\mathbf{x}^{(i)})$ and $\phi_{W}(\mathbf{x}^{(j)})$ respectively. As seen by a deep network, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{W}(\mathbf{x}^{(i)})$ change would change $\phi_{W}(\mathbf{x}^{(j)})$ as well. If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have large effect on $\phi_{W}(\mathbf{x}^{(j)})$. Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have small effect on $\phi_{W}(\mathbf{x}^{(j)})$. To measure the similarity between data points, we employ a symmetric kernel proposed by Charpiat et al. (2019b): The Inner Product
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{W}\phi_{W}(\mathbf{x}^{(i)}),\nabla_{W}\phi_{W}(\mathbf{x}^{(j)})\right\rangle,$ (1)
@@
We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(W\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$. We have $\phi_{W}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. For simplicity, we remove the bias term. Let denote $\mathbf{u}=W\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. Using the chain rule:
$\nabla_{W}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial W}\right)$
$=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial W}\right)$
@@
[…]
@@
The first term on the right-hand side of Eqn. 3 is the partial derivatives of the loss w.r.t the predicted output $\hat{\mathbf{y}}$. Regarding to the fact: $\mathbf{y}$ is the one-hot vector present label $k$ has element $y_{k}=1$ and $y_{i}=0$ if $i\neq k$. We have:
@@
The second term is the matrix comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t $\mathbf{u}$. Regarding to the fact:
@@
Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{k}$:
@@
Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{i}$:
@@
Using the chain rule, we get:
@@
Substitute Eqn. 7 and Eqn. 9 into Eqn. 6, we get a column vector:
@@
The inner product kernel in Eqn. 2 become:
@@
We denote $\mathbf{G}^{(ij)}$ as the dot product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$. Suppose input $\mathbf{x}^{(i)}$ and input $\mathbf{x}^{(j)}$ have label $k$ and $k^{\prime}$ corresponding. We have
@@
During the training process, the model is more confident about the labels of data points, indicating the value of $\hat{y}^{(i)}_{k}$ and $\hat{y}^{(j)}_{k^{\prime}}$ being closer to $1$. Assume a well-trained model, and $\hat{y}^{(i)}_{k}\approx\hat{y}^{(j)}_{k^{\prime}}=\alpha$; $\hat{y}^{(i)}_{n}=\hat{y}^{(j)}_{n}\approx\epsilon=\frac{1-\alpha}{N-1}$. ($n\neq k$ and $n\neq k^{\prime}$). Substitute these values into Eqn. 13 and Eqn. 14, we get :
@@ 3.3 Algorithm
Error Detection.
Given a noisy dataset $\mathcal{D}$, for each data point $\mathbf{z}^{(i)}\in\mathcal{D}$ with label $\mathbf{y}^{(i)}$, our algorithm finds $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ such that every data point in $\mathcal{D}$ but not in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ is at most similar to $\mathbf{z}^{(i)}$ as the least similar point in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 8). We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9). The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds. The lower $\mathbf{s}$ is, the more likely a label error. We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11).
Error Rectification.
We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$. Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
Algorithm 1 Error Detection and Rectification
@@
4 Experiment
4.1 Experiment Setting
(a) Error detection accuracy of methods measure on Snippets. The x-axis in the figures presents the change of $t$ from $10\to 100\%$.
Table 1: Test accuracy after removing/rectifying potential noise samples on Snippets with 20% noise. The best and runner-up are marked.

### v2

@@ Abstract
Label noise in datasets could significantly damage the performance and robustness of deep neural networks (DNNs) trained on these datasets. As the size of modern DNNs grows, there is a growing demand for automated tools for detecting such errors. In this paper, we propose post-hoc, model-agnostic noise detection and rectification methods utilizing the penultimate feature from a DNN. Our idea is based on the observation that the similarity between the penultimate feature of a mislabeled data point and its true class data points is higher than that for data points from other classes, making the probability of label occurrence within a tight, similar cluster informative for detecting and rectifying errors. Through theoretical and empirical analyses, we demonstrate that our approach achieves high detection performance across diverse, realistic noise scenarios and can automatically rectify these errors to improve dataset quality. Our implementation is available at https://anonymous.4open.science/r/noise-detection-and-rectification-AD8E.
Keywords: Noisy label detection and rectification, post-hoc, model-agnostic, DNN, LLM
[…]
Detecting and Rectifying Noisy Labels:
A Similarity-Based Approach
Dang Huu-Tien, Minh-Phuong Nguyen, and Naoya Inoue†,‡
JAIST,  RIKEN
{tiendh, phuongnm, naoya-i}@jaist.ac.jp
Abstract content
1.  Introduction
@@
While the majority of knowledge in Artificial Intelligence (AI) systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f. the LeCun’s cake). For LLMs such as GPTs Brown et al. (2020b), Llama Touvron et al. (2023a, b), and Gemini Team et al. (2023), supervised learning accounts for only a small fraction of the total computation budget but has a significant impact on the models’ performance.
The efficacy and resilience of supervised models are susceptible to degradation due to noise in datasets. Noise (mislabeled) samples could result from mistakes in the labeling process or human disagreement Goh and Mueller (2023); Goh et al. (2022). Human error rates on real-world datasets could range from a few to tens of percent Beyer et al. (2020); Northcutt et al. (2021b). For more challenging datasets where expert knowledge is required, the error rate could be even higher. Recent research (e.g., Zhou et al., 2023; Gunasekar et al., 2023) finds that high-quality training data significantly improves performance while reducing the training cost by orders of magnitude. The large scale of modern datasets makes manual data inspection prohibitively expensive. The immense complexity of deep models makes it extremely hard to reliably and efficiently predict the effect a data point has on the performance of a deep model Basu et al. (2021).
The need for automated tools for improving the quality of supervised learning data is rising as datasets and LLMs are getting larger at an unprecedented speed. Previous works Dau et al. (2022); Nguyen-Duc et al. (2023) showed that removing noise from the training set improves the performance of AI models trained on that dataset. Automatic label noise rectification, however, is an underexplored topic. In this paper, we make the following contributions:
  1. (1)
    We first provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Section 3.2).
  2. (2)
    Inspired by this observation, we develop a simple yet effective similarity-based approach with two representative methods for detecting and rectifying label noise in large-scale datasets (Section 3.3). Our methods are posthoc and model-agnostic, that is, they can be applied to any DNN architectures.
  3. (3)
    Extensive experiments demonstrate the superior performance of the proposed methods over gradient-based and confidence-based approaches, across diverse noise scenarios and scales (Section 5).
2.  Background and Related Work
@@
2.1.  Confidence-Based Approach
@@
Confidence-based methods are based on the notion of confident learning Northcutt et al. (2021a) that derives label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al. (2022). Low confidence serves as a heuristic indicating the likelihood of label noise. Given a data point $\mathbf{z}$ with output label $\mathbf{y}=(y_{1},...,y_{k},...y_{N})$, the model’s predicted probabilities is $\mathbf{p}=(p_{1},...,p_{k},...,p_{N})$ over $N$ classes. Northcutt et al. (2021a) proposed three label quality scoring methods:
@@
Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$:
$\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}};\;k\in\{1,2,...,N\}$ (1)
@@
Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class:
$\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}};\;j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$ (2)
Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy:
$\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{S_{N}(\mathbf{p})};S_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$ (3)
$2.$  Gradient-Based Approach
@@
Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN without retraining. Several IF variants with faster measures have been proposed. Dau et al. (2022) proposed a way to adapt IF and its variants, that is, Gradient Dot Product (GD; Charpiat et al. (2019a)), Gradient Cosine (GC; Charpiat et al. (2019a)), and Tracing Gradient Decent (TracIn; Pruthi et al. (2020)), for identifying errors in large-scale source code datasets. The idea is that the gradients of noisy data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points. The algorithm computes the influence score of each data point in the noisy dataset with data points in a reference set. A more negative influence score means it is more likely to be a noisy data point. Nguyen-Duc et al. (2023) use class information to improve the performance and stability of these gradient methods. We summarize the gradient-based approach as follows:
$\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=-\frac{1}{n}\bm{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\bm{H}_{\hat{\bm{\theta}}}^{-1}\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(j)})$ (4)
$\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=\left\langle\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(i)}),\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(j)})\right\rangle$ (5)
[…]
@@
$\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=\sum_{t=1}^{T}\eta_{t}\left\langle\bm{g}_{\bm{\theta}^{(t)}}(\mathbf{z}^{(i)}),\bm{g}_{{\bm{\theta}^{(t)}}}(\mathbf{z}^{(j)})\right\rangle$ (7)
where $\bm{H}^{-1}_{\hat{\bm{\theta}}}$ is the inverse hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ and $\bm{\theta}^{(t)}$ are the learning rate and optimal weights, at checkpoint $t$-th, respectively. $||\cdot||$ denote the Euclidean norm, $\langle\cdot,\cdot\rangle$ denote the inner product.
2.3.  Other Approaches
The rule-based approach Chu et al. (2013) and statistics-based approach Huang and He (2018) are commonly used for structured data such as tabular data. These methods are not suitable for deep learning, as they assume convexity in the model, and the rules in many large-scale datasets are not easy to find and describe. Bahri et al. (2020) use kkNN on the logit of a DNN to filter out noisy samples. Gao et al. (2016); Reeve and Kabán (2019) focus on examining the resilience of kkNN to random noise. Wu et al. (2020) proposed TopoFilter that filters noisy by utilizing the kkNN and Euclidean distance between pre-logits. Chong et al. (2022) uses the negative log probability of the assigned classes, akin to the Self-Confidence method. NoiseGPT Wang et al. (2024) leverages Multimodal LLMs to measure the “probability curvature” of data points. This curvature differs between benign and noisy samples. NoiseGPT then utilizes the In-Context Discrepancy to suggest or select corrected labels, serving as a knowledge expert for label noise detection and rectification. Unlike these prior works, this work proposes a post hoc, model-agnostic similarity-based approach for both detection and rectification, with theoretical justification under diverse, human-oriented noise.
3.  Methodology
3.1.  Motivation
As a first start, we design a control experiment that randomly selects data points from the dataset and replaces their labels with random labels sampled from the output space.
@@ 3.1.  Motivation
(a) Distribution of (a) Cosine similarity and (b) Dot product over IMDB Maas et al. (2011) with $10\%$ noise. Blue bars represent the similarity between mislabeled data points and their true class data points, red bars represent the similarity between mislabeled data points and other class data points. Features are obtained from a trained BERT.
We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points in the penultimate latent space, i.e., the representation space before softmax. We then employ the cosine similarity and the dot product similarity to compute the similarity between the mislabeled data points and data points from other classes. Surprisingly, as an illustrative example in Fig. 2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points. We hypothesize that in overparameterized models, when the noise rate is sufficiently small, around $5-10\%$—which is a realistic assumption in practice—the model tends to memorize the noise while approaching the optimum, causing the gradient contribution of the noisy samples to diminish. Consequently, as the model converges, the influence of noise on the decision boundary becomes negligible. To support this observation, we further provide a theoretical analysis in Section 3.2.
3.2.  Theoretical Analysis
@@
Let us consider a DNN that receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension. Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias term. Given a function $\phi_{\bm{W}}$ parameterized by $\bm{W}\in\mathbb{R}^{N\times d}$. For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we can obtain two output vectors $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ and $\phi_{\bm{W}}(\mathbf{x}^{(j)})$ from $\phi_{\bm{W}}$, respectively. As seen by a DNN, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ change would change $\phi_{\bm{W}}(\mathbf{x}^{(j)})$ as well. If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ have large effect on $\phi_{\bm{W}}(\mathbf{x}^{(j)})$. Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ have small effect on $\phi_{\bm{W}}(\mathbf{x}^{(j)})$. To measure the similarity between data points, we employ a symmetric kernel proposed by Charpiat et al. (2019b): The Inner Product kernel
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{\bm{W}}\phi_{\bm{W}}(\mathbf{x}^{(i)}),\nabla_{\bm{W}}\phi_{\bm{W}}(\mathbf{x}^{(j)})\right\rangle,$ (8)
@@
We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(\bm{W}\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$. We have $\phi_{\bm{W}}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};\bm{W})$. For simplicity, we remove the bias term. Let denote $\mathbf{u}=\bm{W}\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. Using the chain rule, we obtain:
$\nabla_{\bm{W}}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\bm{W}}\right]=\textnormal{vec}\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial\bm{W}}\right]$
@@
where vec[$\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\bm{W}}\right]$ is the vectorization of the derivative of the loss $\ell$ w.r.t. $\bm{W}$. The partial $\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}$ is
@@
The first term on the right-hand side of Eqn. 10 is the partial derivatives of the loss w.r.t. the predicted output $\hat{\mathbf{y}}$. Regarding to the fact: $\mathbf{y}$ is the one-hot vector present label $k$ has element $y_{k}=1$ and $y_{i}=0$ if $i\neq k$. We have:
@@
The second term is the matrix that comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t. $\mathbf{u}$. Regarding the fact:
@@
Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t. $u_{k}$:
@@
Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t. $u_{i}$:
@@
Using the chain rule, we obtain:
@@
Substitute Eqn. 14 and Eqn. 16 into Eqn. 13, we get a column vector:
@@
The inner product kernel in Eqn. 8 becomes:
@@
We denote $\mathbf{G}^{(ij)}$ as the product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$. Suppose input $\mathbf{x}^{(i)}$ and input $\mathbf{x}^{(j)}$ have label $k$ and $k^{\prime}$ corresponding. We have
@@
During the training process, the model is more confident about the labels of data points, indicating the value of $\hat{y}^{(i)}_{k}$ and $\hat{y}^{(j)}_{k^{\prime}}$ being closer to $1$. Assume a well-trained model, and $\hat{y}^{(i)}_{k}\approx\hat{y}^{(j)}_{k^{\prime}}=\alpha$; $\hat{y}^{(i)}_{n}=\hat{y}^{(j)}_{n}\approx\epsilon=\frac{1-\alpha}{N-1}$. ($n\neq k$ and $n\neq k^{\prime}$). Substitute these values into Eqn. 20 and Eqn. 21, we get:
@@
Algorithm 1 Noise Detection and Rectification
@@
We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9). The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds. The lower $\mathbf{s}$ is, the more likely a label noise. We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11). We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$. Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
4.  Experiment
4.1.  Dataset and Model
We use Snippets Phan et al. (2008) and IMDB Maas et al. (2011) for evaluation. Snippets is a dataset of web search snippets retrieved from Google Search with 8 domains, including business, computers, culture-arts-entertainment, education-science, engineering, health, politics-society, and sports. The training data and testing data include $10,060$ and $2,280$ snippets, respectively. For validation purposes, we randomly split the original training data into train/validation with the ratio of $8048/2012$. IMDB is the most common benchmark for sentiment analysis tasks. IMDB includes $50,000$ reviews from the Internet Movie Database website with original $25,000$ negative and $25,000$ positive reviews. For validation purposes, we randomly split into training, validation, and test sets of sizes $15,000$, $5,000$, and $25,000$. We use BERT Devlin et al. (2019) base uncased version, which is one of the most standard used models for language understanding tasks. We use open-weight LLMs: Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B AI@Meta (2024) for the zero-shot detection scenarios in Subsection 5.2. Due to space constraints, we present the main results and include some additional results in the Appendix upon the final decision.
4.2.  Experimental Setup
(a) Noise detection accuracy of methods measured on Snippets (Fig. 8(a)a-c) and IMDB (Fig. 8(a)d-f) across various sizes and noise types.
Table 1: Test accuracy of model measured on Snippets after removing or rectifying top $20\%$ potential noisy samples ranked by methods. The best, improvement, drop, and runner-up are marked.